OpenStudy (anonymous):

Kim earns $30 for babysitting on Friday nights. She makes an average of $1.25 in tips per hour. Write the function of Kim's earnings, and solve for how much she would make after 3 hours. y = 1.25x + 30; $33.75 y = 30x + 1.25; $43.75 y = 30x + 1.25; $91.25 y = 1.25x + 30; $31.25

OpenStudy (neer2890):

Kim earns $30 for babysitting on Friday nights. so if she's doesn't get any tip then also at the end of night, she has $30. but she earns $1.25 per hour. Let x is the number of hours Kim got tips. so total earnings of Kim in x hours =1.25*x Total earnings of x hours and babysitting= 1.25x+30 so, in 3 hours, total earnings of Kim = 1.25*3+30= $33.75